The volume of 1.204×10^24 molecules of water at 4 cecius is

Grade:11

1 Answers

Akm
5 years ago
Volume of 1 mole water= 1000g/l
1 mole contains avogardos no of molecules.
So volume of 1.204×10^24 molecule=( 1000÷(6.023×10^23))×1.204×10^24
=1.37×10^-45